OpenStudy (anonymous):

Multiple and simplify by factoring √6 . √21

OpenStudy (asnaseer):

The way to do these is to split each number into its prime factors. so: 6 = 2*3 21 = 3*7 Then you can do:\[\sqrt{6}*\sqrt{21}=\sqrt{6*21}=\sqrt{2*3*3*7}=\sqrt{3^2*2*7}=3\sqrt{2*7}=3\sqrt{14}\]